# Learning Naturally Aggregated Appearance for Efficient 3D Editing

### v1

@@ Abstract
Neural radiance fields, which represent a 3D scene as a color field and a density field, have demonstrated great progress in novel view synthesis yet are unfavorable for editing due to the implicitness. In view of such a deficiency, we propose to replace the color field with an explicit 2D appearance aggregation, also called canonical image, with which users can easily customize their 3D editing via 2D image processing. To avoid the distortion effect and facilitate convenient editing, we complement the canonical image with a projection field that maps 3D points onto 2D pixels for texture lookup. This field is carefully initialized with a pseudo canonical camera model and optimized with offset regularity to ensure naturalness of the aggregated appearance. Extensive experimental results on three datasets suggest that our representation, dubbed AGAP, well supports various ways of 3D editing (e.g., stylization, interactive drawing, and content extraction) with no need of re-optimization for each case, demonstrating its generalizability and efficiency. Project page is available at https://felixcheng97.github.io/AGAP/.
Figure 1: AGAP aggregates 3D appearance as natural 2D canonical images. With 2D image processing tools, our method enables various ways of 3D editing, including (a) scene stylization, (b) content extraction, and (c) interactive drawing, without further re-optimization.
@@ 1 Introduction
While recent advancements in 3D representations like neural radiance fields (NeRF) [31] have shown impressive reconstruction capabilities for real-world scenes, the need for further progress in 3D editing arises as the desire to recreate and manipulate these scenes. The field of 3D editing has witnessed significant development in recent years. Traditional 3D modeling approaches [21, 43, 42, 55] typically rely on reconstructing scenes using meshes. By combining meshes with texture maps, we can enable appearance editing during the rendering process. However, these methods usually face difficulties in obtaining detailed and regular texture maps, typically in complex scenes, which significantly hinders effective editing and compromises user-friendliness.
@@
Recent neural radiance fields offer high-quality scene reconstruction capabilities, but manipulating the implicit 3D representation embedded within neural networks is inherently complex and non-straightforward. Existing NeRF-based editing approaches can be mainly divided into two categories: some methods like [8, 9, 52, 56, 59, 63] target geometry editing, usually taking advantage of meshes, while the other [14, 36, 24, 60, 62] focuses on 3D stylization using images or texts as style guidance. However, re-optimizing the pre-trained NeRF models is necessary to incorporate the desired editing effects into the underlying 3D representation, resulting in time-consuming processes. Consequently, it is crucial to develop a user-friendly framework that can efficiently and effectively support various edits within a single model.
@@
In this paper, we introduce a novel editing-friendly representation AGAP with naturally Aggregated Appearance, consisting of a 3D density grid for geometry estimation and a canonical image plus a projection field for appearance or texture modeling. Our method attempts to link the 3D representation with a natural 2D aggregated canonical representation. Concretely, a learnable canonical image is designed as the interface for editing, which aggregates the appearance by projecting the 3D radiance to a natural-looking image by the associated projection field. To ensure the naturalness of the aggregated canonical image with strong representation capacity, the associated projection field is carefully initialized by a pseudo canonical camera model and complemented by a learned view-dependent offset. The underlying 3D structure is modeled by an explicit 3D density grid. After a scene is optimized through volume rendering, AGAP supports various ways of 3D editing in a user-friendly setting by applying different 2D image processing tools on the canonical images without re-optimization of the original model. We evaluate the effectiveness of our method on three datasets, including LLFF [30], Replica [46, 13], and Instruct-NeRF2NeRF [14] in various editing tasks, which are scene stylization, content extraction, and texture editing (i.e., drawing). Experimental results further show that AGAP can achieve on-par performance with NeRF [31] in terms of PSNR with a hash grid-based projection field.
@@ 2 Related Work
Neural scene editing. Existing research on NeRF editing can be broadly categorized into two: one focuses on editing the geometry [8, 9, 52, 56, 59, 63]; the other, known as style-based editing [14, 36, 24, 60, 62], aims to achieve scene stylization. Our research aligns with the latter category. Many NeRF stylization methods [60, 16, 33] have adopted techniques from 2D image stylization with style loss and content loss on images for NeRF optimization. While these methods can deliver 3D-consistent editing, they are primarily limited to global texture modifications and lack flexibility. Later, CLIP-NeRF [49] incorporates text conditions by regularizing the CLIP embeddings of the global scene with input prompts. Subsequent studies [24] extract 2D features such as DINO [5, 34] for local editing. Most recently, Instruct-Nerf2Nerf [14] proposes an iterative approach to edit the input images using pre-trained diffusion models [4] for underlying NeRF optimization. Despite achieving high-fidelity editing results, these methods necessitate optimization for each text prompt or reference image, which is inefficient.
Neural atlases. Our work shares similar insights with the research area of neural atlases [20, 58, 35], which decompose videos into a canonical form with learned deformations, thereby enabling consistent video editing. Approaches such as neural layered atlases [20] employ an implicit network to distinguish foreground and background movement, dividing them into distinct layers. The CoDeF [35] methodology represents 2D videos using content deformation fields by integrating 3D hash tables [32]. However, these approaches lack 3D priors, limiting the effectiveness of 3D viewpoint changes in 3D scene editing.
Figure 2: The overall pipeline. AGAP consists of two components: (1) an explicit 3D density grid $\phi_{G}$ to estimate geometry for density $\sigma$; (2) an explicit canonical image $\phi_{I}$ with an associated view-dependent projection field $P$ to aggregate appearance for color $\mathbf{c}$. By performing 2D image processing on the canonical image, our method enables various editing (e.g., content extraction, interactive drawing, and scene stylization) through volume rendering without the need for re-optimization.
@@ 3 Method
Formally, given a set of multi-view training images $\mathcal{I}$, our method models the scene appearance by an explicit canonical image $\phi_{I}$ plus a corresponding implicit projection field $P$ inspired by [35, 39, 38]; the scene geometry is estimated by an explicit 3D density grid $\phi_{G}$. With such a representation, one can render different views of the scene through volume rendering. Our key property is that by explicitly editing the canonical image $\phi_{I}$, it can propagate the edited appearance to the whole scene through the projection field $P$ without any further re-optimization. An overview of our framework is shown in Fig. 2.
@@ 3.2 Model Formulation
Our representation disentangles the density and color of the scene and uses different modalities for modeling.
Density. Similar to DVGO [47], we adopt a voxel-grid representation to model the 3D density for an efficient query. Given a particular query point $\mathbf{p}_{xyz}\triangleq\mathbf{r}(t_{i})\in\mathbb{R}^{3}$, we can obtain the corresponding density $\sigma\in\mathbb{R}$ via a trilinear interpolation, followed by a Softplus activation:
@@
where $\phi_{G}$ denotes the one-channel voxel grid with learnable parameter $\phi_{g}$ at a voxel resolution size of $N_{x}\times N_{y}\times N_{z}$. Concerning the fact that in 3D reconstruction, textures are applied to the surface of the mesh to provide visual details such as colors, patterns, and material properties [51, 17], we hope our model can obtain a coarsely accurate density estimation to perceive the 3D geometry at the early training stage, facilitating the subsequent learning of appearance aggregation. Hence, we opt for an explicit voxel-grid representation to achieve fast convergence instead of utilizing an implicit MLP such as NeRF [31]. Such a choice is also proven to be crucial by our experiments in Sec. 4.4.
Appearance. In order to empower explicit 3D editing capabilities, we formulate the color appearance by an explicit canonical image $\phi_{I}\in\mathbb{R}^{H\times W\times 3}$ plus an associated view-dependent implicit projection field $P(\cdot,\cdot):(\mathbb{R}^{3},\mathbb{R}^{3})\to\mathbb{R}^{2}$, where $H$ and $W$ represent the image height and width, respectively. This formulation maps a given query point $\mathbf{p}_{xyz}$ in the 3D field with viewing direction $\mathbf{d}$ to the projected 2D point $\mathbf{p}_{uv}$ on the canonical image $\phi_{I}$.
The design of the projection field is crucial to achieving naturalness and completeness in the learned canonical image. We model the projection field $P$ in a residual manner, which consists of a non-learnable canonical projection initialization $P_{c}$ and a projection offset learning $P_{o}$. The canonical projection initialization $P_{c}$ plays a significant role in ensuring a natural-looking canonical image, while the projection offset $P_{o}$ aims to address view-dependent effects and handle occlusions present in complex scenes.
Specifically, the canonical projection initialization $P_{c}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{2}$ projects the query 3D point $\mathbf{p}_{xyz},$ to an initial 2D projection point $\tilde{\mathbf{p}}_{uv}$ on the canonical image. Further elaboration on the selection of an appropriate projection initializing function $P_{c}$ will be provided in Sec. 3.3. The projection offset $\Delta\mathbf{p}_{uv}$ is modeled by $P_{o}$ with parameter weights $\phi_{P}$:
@@
For simplicity, we omit the 3D positional encoding $\gamma_{p}$ of $\mathbf{p}_{xyz}$ and the viewing direction encoding $\gamma_{d}$ of $\mathbf{d}$ in Eq. 4. The final projection point $\mathbf{p}_{uv}$ can be derived as follows:
$\mathbf{p}_{uv}=\tilde{\mathbf{p}}_{uv}+\Delta\mathbf{p}_{uv}=P_{c}(\mathbf{p}_{xyz})+P_{o}(\mathbf{p}_{xyz},\mathbf{d};\phi_{P}).$ (5)
The projection point $\mathbf{p}_{uv}$ is then used to query the RGB color $\mathbf{c}\in\mathbb{R}^{3}$ from the canonical image $\phi_{I}$ via interpolation:
3.3 Canonical Projection Initialization
Learning a good projection field $P$ for 3D-to-2D projection is indeed a non-trivial research challenge [10], particularly for 360-degree inward-facing scenes [3, 31]. Thus, we relax the 3D scene of interest in our approach, focusing on forward-facing scenes [30] and outward-facing panorama data [13]. The crucial step in achieving a natural-looking canonical image as the aggregated appearance is to position a pseudo canonical camera into the scene for a careful canonical projection initialization as $P_{c}$.
Forward-facing data and NDC representation. Similar to NeRF, we utilize the normalized device coordinate (NDC) space to model the real forward-facing captures. In perspective projection, the 3D points in the truncated pyramid frustum in camera (eye) coordinates are mapped to a cube $[-1,1]^{3}$ in NDC space. The pseudo canonical camera is defined as the average camera pose positioned at the world origin. By doing so, the perspective projection $f_{p}$ can map any given point $\mathbf{p}_{xyz}$ in the world coordinates to its corresponding NDC point $(p_{x^{\prime}},p_{y^{\prime}},p_{z^{\prime}})\triangleq f_{p}(\mathbf{p}_{xyz})$. The coordinate range of the voxel grid $\phi_{G}$ is defined by the bounding box in NDC space. Please refer to the original NeRF paper [31] for the detailed derivation of NDC.
One notable property of NDC is that for a ray $\mathbf{r}$ origin from the camera center, the resulting projected camera ray $\mathbf{r}^{\prime}$ in NDC space is always perpendicular to the $x^{\prime}y^{\prime}$ plane. This implies that all points along $\mathbf{r}^{\prime}$ share identical values for $p_{x^{\prime}}$ and $p_{y^{\prime}}$. Leveraging this property, we can have a simple and appropriate initialization for the projection field:
$\tilde{p}_{u}=p_{x^{\prime}},\quad\tilde{p}_{v}=p_{y^{\prime}},$ (7)
where we have the notation $(\tilde{p}_{u},\tilde{p}_{v})\triangleq\tilde{\mathbf{p}}_{uv}$.
@@ 3.3 Canonical Projection Initialization
Figure 3: Visual comparison of novel-view stylization results on the LLFF dataset given different reference images. Our method can better preserve textural consistencies as highlighted in the figure.
Panorama data and contracted representation. The training images for panorama scenes are typically captured by cameras positioned around the global origin, with outward-facing views covering a 360-degree field of view. We choose to place the pseudo canonical camera at the global origin and define the Equirectangular projected image as a feasible canonical projection initialization.
Drawing inspiration from the smooth coordinate transforms in [3, 40] for 360-degree inward-facing unbounded data, we introduce a new contracted formulation $f_{c}$ specifically for outward-facing (unbounded) panorama scenes:

### v2

@@ Abstract
Neural radiance fields, which represent a 3D scene as a color field and a density field, have demonstrated great progress in novel view synthesis yet are unfavorable for editing due to the implicitness. This work studies the task of efficient 3D editing, where we focus on editing speed and user interactivity. To this end, we propose to learn the color field as an explicit 2D appearance aggregation, also called canonical image, with which users can easily customize their 3D editing via 2D image processing. We complement the canonical image with a projection field that maps 3D points onto 2D pixels for texture query. This field is initialized with a pseudo canonical camera model and optimized with offset regularity to ensure the naturalness of the canonical image. Extensive experiments on different datasets suggest that our representation, dubbed AGAP, well supports various ways of 3D editing (e.g., stylization, instance segmentation, and interactive drawing). Our approach demonstrates remarkable efficiency by being at least 20$\times$ faster per edit compared to existing NeRF-based editing methods. Project page is available at https://felixcheng97.github.io/AGAP/.
Figure 1: AGAP aggregates 3D appearance as natural 2D canonical images. With image processing tools, AGAP enables various ways of 3D editing without re-optimization, including (a) scene stylization, (b) instance segmentation, and (c) interactive drawing.
Table 1: AGAP supports various editing cases. Our editing time per edit is significantly shorter thanks to the optimization-free editing pipeline. Note that the times listed below do not include the comparable pre-training and rendering times across methods.
[…]
ARF [84] ✓ ✗ ✗ ✗ 371s
Ref-NPR [87] ✓ ✗ ✗ ✗ 514s
DFFs [35] ✗ ✓ ✓ ✗ 516s
IN2N [$\bm{-}$ [^1] Possible but highly depends on prompts ✗ ✗ $\sim$10000s
GaussianEditor [10] ✓ ✓ $\bm{-}$ [^2] Not inherently support in the official implementation ✗ 1320s
Ours ✓ ✓ ✓ ✓ 20s
@@ 1 Introduction
While recent advancements in 3D representations like neural radiance fields (NeRF) [46] have shown impressive reconstruction capabilities for real-world scenes, the need for further progress in 3D editing arises as the desire to recreate and manipulate these scenes. The field of 3D editing has witnessed significant development in recent years. Traditional 3D modeling approaches [30, 62, 61, 77] typically rely on reconstructing scenes using meshes. By combining meshes with texture maps, we can enable appearance editing during the rendering process. However, these methods usually face difficulties in obtaining detailed and regular texture maps, typically in complex scenes, hindering effective editing and user-friendliness.
@@
Recent neural radiance fields offer high-quality scene reconstructions, but manipulating the implicit 3D representation embedded within neural networks is inherently non-straightforward. Existing NeRF-based editing approaches can be mainly divided into two categories: some methods like [9, 11, 71, 79, 83, 86] target geometry editing, usually taking advantage of meshes, while the other [18, 52, 35, 84, 87] focuses on 3D stylization using images or texts as guidance. However, re-optimizing the original NeRF models is necessary to incorporate the desired editing effects into the underlying 3D representation, resulting in time-consuming processes. Consequently, it is crucial to develop a user-friendly framework that can efficiently and effectively support various edits within a single model.
@@
This paper introduces a novel editing-friendly representation AGAP with naturally Aggregated Appearance for efficient 3D editing, consisting of a 3D density grid for geometry estimation and a canonical image plus a projection field for appearance modeling. Our method attempts to link the 3D representation with a natural 2D canonical representation. Concretely, a learnable canonical image is designed as the interface for editing, which aggregates the appearance by projecting the 3D radiance to a natural-looking image by the associated projection field. To ensure the naturalness of the aggregated canonical image with strong representation capacity, the projection field is carefully initialized using a pseudo canonical camera model and complemented by a learned view-dependent offset. The underlying 3D structure is modeled by an explicit 3D density grid.
AGAP supports various ways edits of a 3D scene in a user-friendly and efficient manner by applying different 2D image processing tools on the canonical images without re-optimization. Tab. 1 summarizes the comparison of our method with existing methods in terms of editing functionalities and per-edit efficiency. We evaluate the effectiveness on various datasets, including LLFF [45], DTU [24], NeUVF [41], Replica [65, 17], IN2N [18], NeRF-Synthetic [46], and Mip-NeRF 360 [3] datasets in various editing tasks which are scene stylization, instance segmentation, and texture editing (i.e., drawing). Experimental results show AGAP support various 3D editing tasks with on-par performance but at least 20$\times$ faster per edit.
@@ 2 Related Work
Neural scene editing. Existing research on NeRF editing can be broadly categorized into two: one focuses on editing the geometry [9, 11, 71, 79, 83, 86]; the other, known as style-based editing [18, 52, 35, 84, 87], aims to achieve scene stylization. Our research aligns with the latter category. Many NeRF stylization methods [84, 22, 49] have adopted techniques from 2D image stylization with style loss and content loss on images for NeRF optimization. While these methods can deliver 3D-consistent editing, they are primarily limited to global texture modifications and lack flexibility. Later, CLIP-NeRF [68] incorporates text conditions by regularizing the CLIP embeddings of the global scene with input prompts. Subsequent studies [35] extract 2D features such as DINO [5, 50] for local editing; IN2N [18] proposes an iterative approach to edit the input images using pre-trained diffusion models [4] for underlying NeRF optimization. Despite achieving high-fidelity editing results, most NeRF-based methods [8, 13, 19, 20, 26, 32, 36, 40, 43, 47, 57, 59] necessitate optimization for each text prompt or reference image, which can be inefficient in practical use. More recently, many works [10, 14, 56, 73, 74, 76, 78, 81, 23] start explore 3D editing with Gaussian Splatting [31]. , As we generally focus on NeRF-based editing alternatives, we choose GaussianEditor [10] among them as the representative comparing baseline.
Neural atlases. Our work shares similar insights with the research area of neural atlases [29, 82, 51, 41], which decompose videos into a canonical form with learned deformations, thereby enabling consistent video editing. Approaches such as neural layered atlases [29] employ an implicit network to distinguish foreground and background movement, dividing them into distinct layers. The CoDeF [51] methodology represents 2D videos using content deformation fields by integrating 3D hash tables [48]. However, these approaches lack 3D priors, limiting the effectiveness of 3D viewpoint changes in 3D scene editing.
Figure 2: The overall pipeline. AGAP consists of two components: (1) an explicit 3D density grid $\phi_{G}$ to estimate geometry for density $\sigma$; (2) an explicit canonical image $\phi_{I}$ with an associated view-dependent projection field $P$ to aggregate appearance for color $\mathbf{c}$. By performing 2D image processing on the canonical image, our method enables various editing (e.g., instance segmentation, interactive drawing, and scene stylization) through volume rendering without the need for re-optimization.
@@ 3 Method
As shown in Tab. 1, existing editing methods [84, 87, 18, 10] necessitate several minutes or even hours per edit to re-optimize their original models. Meanwhile, we argue that their implicit editing procedures through re-optimization reduce the level of user interactivity compared to explicit editing ways. In view of such deficiencies, the core concept of AGAP is to learn a 2D canonical image as the interactive medium and allow users to do efficient and explicit 3D editing by lifting image processing on the 2D canonical image. Under such an idea, the key challenges involve designing a projection field that bridges the appearance of the 3D scene appearance with the 2D canonical image and ensuring the naturalness of the canonical image.
@@ 3.2 Model Formulation
Formally, given a set of multi-view training images $\mathcal{I}$, AGAP models the scene appearance by an explicit canonical image $\phi_{I}$ plus a corresponding implicit projection field $P$ inspired by [51, 55, 54]; the scene geometry is estimated by an explicit 3D density grid $\phi_{G}$. With such a representation, one can render different views through volume rendering. Our key property is that by explicitly editing the canonical image $\phi_{I}$, it can propagate the edited appearance to the whole scene through the projection field $P$ without any re-optimization. Our overall framework is shown in Fig. 2.
Density. In 3D modeling, textures are applied to the mesh surface to provide visual details such as colors, patterns, and material properties [70, 25]. We opt for an explicit voxel-grid representation [66, 15, 7, 48] instead of an implicit MLP such as NeRF [46] to achieve fast convergence and efficient query. Given a particular query point $\mathbf{p}_{xyz}\triangleq\mathbf{r}(t_{i})\in\mathbb{R}^{3}$, we obtain the corresponding density $\sigma\in\mathbb{R}$ via a trilinear interpolation, followed by a Softplus activation:
@@
where $\phi_{G}$ denotes the one-channel voxel grid with learnable parameter $\phi_{g}$ at a voxel resolution size of $N_{x}\times N_{y}\times N_{z}$. With a density grid with explicit parameterization, we hope our model can obtain a coarsely accurate density estimation of the 3D geometry at the early training stage, facilitating the learning of 2D appearance aggregation. Such a choice is also proven to be crucial by our experiments in Sec. 4.4.
Appearance. In order to empower efficient 3D editing capabilities, we formulate the color appearance by an explicit canonical image $\phi_{I}\in\mathbb{R}^{H\times W\times 3}$ with an associated view-dependent implicit projection field $P(\cdot,\cdot):(\mathbb{R}^{3},\mathbb{R}^{3})\to\mathbb{R}^{2}$, where $H$ and $W$ represent the image height and width, respectively. This formulation maps a given query point $\mathbf{p}_{xyz}$ in the 3D field with viewing direction $\mathbf{d}$ to the projected 2D point $\mathbf{p}_{uv}$ on the canonical image $\phi_{I}$. The projection point $\mathbf{p}_{uv}$ is then used to query the RGB color $\mathbf{c}\in\mathbb{R}^{3}$ from the canonical image $\phi_{I}$ via interpolation:
3.3 Canonical Projection with Projection Offset
We model the projection field $P$ as a non-learnable canonical projection $P_{c}$ with a projection offset learning $P_{o}$. To model different 3D scenes, we choose suitable foundational projection as the canonical projection $P_{c}$ derived from different real camera models, which plays a significant role in ensuring naturalness and completeness in the learned canonical. The projection offset $P_{o}$ aims to address view-dependent effects and handle occlusions in complex scenes.
Specifically, the canonical projection $P_{c}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{2}$ projects the query 3D point $\mathbf{p}_{xyz}$ to an initial 2D projection $\tilde{\mathbf{p}}_{uv}=P_{c}(\mathbf{p}_{xyz})$ on the canonical. The projection offset $\Delta\mathbf{p}_{uv}$ is modeled by $P_{o}$ with parameter weights $\phi_{P}$:
@@ 3.3 Canonical Projection with Projection Offset
For simplicity, we omit the 3D positional encoding $\gamma_{p}$ of $\mathbf{p}_{xyz}$ and the viewing direction encoding $\gamma_{d}$ of $\mathbf{d}$ in Eq. 5.
The final projection point $\mathbf{p}_{uv}$ is formulated as follows:
$\mathbf{p}_{uv}=\tilde{\mathbf{p}}_{uv}+\Delta\mathbf{p}_{uv}.$ (6)
3.4 Optimization and Regularization
$\mathcal{L}_{uv}=\left\|\Delta\mathbf{p}_{uv}\right\|_{2}^{2}.$ (7)
Optimization objective. The final optimization process of our method is formulated as follows:
IN2N [18]
GaussianEditor [10]
Ours
Refs
IN2N [18]
GaussianEditor [10]
Ours
@@ 3.4 Optimization and Regularization
Figure 3: Visual comparison of novel-view scene stylization results on the IN2N and LLFF dataset given different text prompts or image reference. Our method can achieve on-par stylization results with the baselines while requiring no time-consuming re-optimization procedures. As highlighted in row two, our method can better preserve color and textural consistencies aligning with the image reference.
Figure 4: By performing explicit edits on the canonical image $\phi_{I}$, our model propagates the editing effects through the learned projection field $P$ for efficient 3D editing.
4 Experiments
Recall that the canonical projection $P_{c}$ can be viewed as positioning a pseudo-canonical camera within the 3D scene. This the crucial step in achieving a natural-looking canonical image $\phi_{I}$ by initializing the projection field $P$, by choosing a suitable foundational projection derived from different real camera models for the canonical projection $P_{c}$ when modeling different types of 3D data scenes.
4.1 Data Types
Forward-facing data. The LLFF [45] dataset comprises real-world forward-facing scenes, where each scene is accompanied by several training images captured by handheld cameras placed in a rough grid pattern nearly on a vertical plane. We evaluate the dataset at a resolution of $756\times 1008$.
We utilize the normalized device coordinate (NDC) space to model the forward-facing captures, where the pseudo canonical camera is defined as the average camera pose at the world origin. The perspective projection $f_{p}$ can map any given point $\mathbf{p}_{xyz}$ in the world coordinates to its corresponding NDC point $f_{p}(\mathbf{p}_{xyz})\triangleq(p_{x^{\prime}},p_{y^{\prime}},p_{z^{\prime}})$. The coordinate range of the voxel grid $\phi_{G}$ is defined by the bounding box in NDC space. One notable property of NDC is that for a ray $\mathbf{r}$ origin from the camera center, all points along the ray share identical values for $p_{x^{\prime}}$ and $p_{y^{\prime}}$. Leveraging this property, we can design an appropriate choice for the canonical projection: $\tilde{\mathbf{p}}_{uv}\triangleq(p_{x^{\prime}},p_{y^{\prime}})$.
Panorama data. The Replica [65] dataset is a collection of various high-quality and high-resolution 3D reconstructions of indoor scenes with clean and dense geometry. We evaluate the panorama scenes as processed in SOMSI [17], where each scene is rendered as a grid of equally spaced spherical images at a resolution of $1024\times 512$.
The panorama data captures outward-facing views covering a 360-degree field of view around the global center. We place the pseudo-canonical camera at the global origin and define the Equirectangular projection as the canonical projection. Drawing inspiration from the smooth coordinate transforms in [3, 58], we introduce a new contracted formulation $f_{c}$ specifically for panorama scenes: